solve the system
7x-y=31
2x+3y=-1

Answer by ikleyn(52798)

.
solve the system

7x - y  = 31,   (1)
2x + 3y = -1.   (2)
~~~~~~~~~~~~~~~~~~~~~~~~


From (1), express y = 31-7x and substitute it into (2). You will get

2x + 3*(31-7x) = -1.

Simplify and solve for x.

It is called "the Substitution method".

On solving linear systems of algebraic equations in two unknowns seee the lessons
    - Solution of a linear system of two equations in two unknowns by the Substitution method
    - Solution of a linear system of two equations in two unknowns by the Elimination method
    - Solution of a linear system of two equations in two unknowns using determinant
    - Geometric interpretation of a linear system of two equations in two unknowns
    - Solving word problems using linear systems of two equations in two unknowns
in this site.